Assuming a company has the below unlevered free cash flow, what is the present value of the terminal value assuming 15.0% discount rate and a 3.0% perpetuity rate? Cash flow for years 1-5 are - $125, $150, $165, $180 and $200.
Answer #1

Please provide rating..

terminal value = 6th year cash flow/(required rate - perpetuity rate)
6th year cash flow = 5th year cash flow *(1+perpetuity rate)
6th year cash flow = =200*103%
6th year cash flow = 206
Therefore terminal value = =206/(15%-3%)
Therefore terminal value =               1,716.67
present value of terminal value = =1716.67/(1.15)^5
present value of terminal value =              853.49